OpenStudy (anonymous):

Let y=(sqrt 6−x) . Find the differential dy when x=3 and dx=0.4 Find the differential dy when x=3 and dx=0.04

OpenStudy (agent0smith):

First find dy/dx \[\large y = \sqrt{6-x} = (6-x)^{1/2}\]

OpenStudy (anonymous):

the derivative is 1/2(6-x)^-1/2(-1)

OpenStudy (agent0smith):

Correct. But write it this way, you'll see why in a moment (i just changed the 1/2 to 0.5 for convenience) \[\large \frac{ dy }{ dx }=-0.5 (6-x)^{-0.5} \] Now lets imagine dy/dx is a fraction...

OpenStudy (agent0smith):

Multiply both sides by dx to get \[\Large dy =-0.5 (6-x)^{-0.5} dx\]

OpenStudy (anonymous):

oh and from there do i plug in the x=3 and the dx right

OpenStudy (agent0smith):

Yes! :D

OpenStudy (anonymous):

alright let me calculate it

OpenStudy (anonymous):

i dont know if i calculate it correctly but i got -1/2sqrt3 for the first one

OpenStudy (agent0smith):

Did you multiply by the dx also? dy when x=3 and dx=0.4 http://www.wolframalpha.com/input/?i=+-1%2F2%286-3%29%5E%28-1%2F2%29*0.4 dy when x=3 and dx=0.04 http://www.wolframalpha.com/input/?i=+-1%2F2%286-3%29%5E%28-1%2F2%29*0.04

OpenStudy (anonymous):

oh i keep forgetting.. alright

OpenStudy (anonymous):

alright i just checked.. thank you hopefully i don't forget later on about the dx

OpenStudy (agent0smith):

You're welcome :)